1. What is a Newtonian fluid? Is water a Newtonian fluid?
2. How does viscousities of (a) liquids and (b) gases vary with
temperature?
1
Expert's answer
2020-08-25T11:13:55-0400
  1. Newtonian liquid is a liquid that is characterized by a linear connection between shear stress and velocity gradient. This makes velocity to be distributed as following:

Aspect ratio between shear stress and velocity gradient is viscosity. Newtonian liquid is a ideal model. In real life it is hard to find truly Newtonian liquids. In most cases, water with a good accuracy can be considered as a Newtonian liquid.

2. (a) liquids: "\mu = A e^{\frac{B}{T}}" , A,B = const.

(b) gases: from elementary calculations using kinetic theory of gases "\mu \propto T^{1/2}"

from more precise theory (Chapman–Enskog theory): "\mu = \mu_0 (\frac{T}{T_0})^{3/2} \frac{T_0+C}{T+C}"